OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. T(13, 2), W( , -5) m = -7/2

OpenStudy (anonymous):

y=m(x-x1)+y1 soo y= -7/2(x-13)+2

OpenStudy (anonymous):

or m= y1-y2/x1-x2 ...........-7/2=2+5/13-x in order to get x by itself multiply both sides by 7 then divide both sides by 13

OpenStudy (anonymous):

3.7?

OpenStudy (anonymous):

no the final answer is -49/26

OpenStudy (anonymous):

-7/2=-3.5 multiply that by 7 to get rid of your fraction on the right then the answer on the left is -24.5 leaving 13x in order to get x by itself divide by on both sides and the final answer is -49/26